# Collective Variable Free Transition Path Sampling with Generative Flow Network

### v2

@@ 4.3 Qualitative results
In this section, we provide a qualitative comparison between our method and baselines and 3D visualizations. For Alanine Dipeptide, we show that our method generates diverse transition paths passing through both transition states. For Polyproline Helix and Chignolin, we visualize the transformation of molecular systems with their potential energy plot. For more visualization, please refer to Appendix C, and project page for animation of the generated transition paths.
Alanine Dipeptide. In Figure 3, we compare transition paths from our method to paths from baselines using the Ramachandran plot. We sample transition paths from $C5$ to $C7_{ax}$ which passes one of the two transition states in the energy landscape indicated by white stars in Figure 3. In this experiment, we sampled 64 transition paths for all cases. The paths from unbiased MD simulations in Figure 3(a) remains trapped in the initial meta-stable region, and thus fail to produce transition paths. Both PIPS [14] and our TPS-GFN generates valid transition paths between $C5$ and $C7_{ax}$ as shown in Figure 3(b) and Figure 3(c), repectively. However, PIPS generates transition paths passing through only one of the transition states while our TPS-GFN generates diverse transition paths, passing through both transition states as shown in Figure 3(c).
Figure 4: Visualization of a transition path from the meta-stable region PP-II to PP-I of Polyproline generated by our TPS-GFN. The top row visualizes the 3d view of states along the time horizon. The bottom row visualizes the potential energy along the time horizon. The backbone of the Polyproline is highlighted in green.
Polyproline Helix and Chignolin. Since the CVs of the Polyproline Helix and the Chignolin are less understood, the Ramachandran plot is unavailable. Alternatively, we visualize the 3D views of the generated transition path at 5 different time steps with their potential energy along the time horizon. In Figure 4, Our method successfully generates transition paths in terms of one CV candidate, handedness [44]. We observe change from the left-handed Helix to the right-handed Helix of the Polyproline and the potential plot which roughly increases up to approximately 440 $\mathrm{fs}$ and decreases after then. Similarly, in Figure 5, our method captures the folding process of the Chignolin since the potential plot which roughly increases up to approximately 950 $\mathrm{fs}$ and decreases after then. We note that our method can produce various lengths of transition paths by cutting paths based on the relaxed indicator function. We also find the average time of the transition from PP-II to PP-I in Polyproline Helix is approximately 720 $\mathrm{fs}$ and the average folding time of chignolin is approximately 1000 $\mathrm{fs}$.
Figure 5: Visualization of a folding process of Chignolin generated by our TPS-GFN. The top row visualizes the 3d view of states along the time horizon. The bottom row visualizes the potential energy along the time horizon.
@@ 5 Conclusion
In this work, we introduced a novel CV-free approach to sampling transition paths using generative flow networks (GFlowNets) to model bias potentials for efficient and diverse sampling. By reformulating the problem as amortized energy-based sampling and employing the one-step trajectory balance (TB) objective, our approach significantly improves the discovery of realistic transition paths compared to existing methods. Evaluations on proteins such as Alanine Dipeptide, Polyproline, and Chignolin demonstrate superior accuracy and diversity in identifying transition states. We believe that our works expand to more complex systems, with significant implications for drug discovery and material design.
@@ Appendix A Generative Flow Networks
In this section, we briefly introduce the concept of GFlowNets and explain how we derive Equation 2. We refer to Lahlou et al. [45] for a comprehensive study on formalizing GFlowNets on continuous state spaces.
Generative flow networks. GFlowNet [15, 16] is a framework to sequentially construct samples $s\in\mathcal{S}$ proportional to a reward function $R(s)$, where $\mathcal{S}$ is a set of data. In other words, GFlowNet aims to learn a target distribution $p^{\ast}(s)$ defined as:
@@
Trajectory balance. For flow matching, Malkin et al. [17] proposed the trajectory balance objective:
@@
where $P_{B}(s_{0}|s_{1};\theta)$ is ignored since $s_{0}$ is the constant initial state. Note that the detailed balance [16] objective also reduces to the same formula as in Equation 5 with one-step GFlowNets.
@@ Appendix B Experiment details
We implement codes based on OpenMM [46] and Pytorch [47] libraries, and run codes on a single GPU (NVIDIA RTX 3090 or RTX A5000).
@@ B.1 OpenMM configurations
We run MD simulations to sample paths and evaluate them at the temperature of 300K, using the Langevin integrator. Other configurations depend on molecules as shown in Table 3.
Table 3: OpenMM configurations depeding on molecules.
Molecule Time horizon Time step Force field Solvent
Alanine Dipeptide 500 $\mathrm{fs}$ 1 $\mathrm{fs}$ amber99sbildn tip3p
Polyproline Helix 10000 $\mathrm{fs}$ 2 $\mathrm{fs}$ amber/protein.ff14SBonlysc implicit/gbn2
Chignolin 5000 $\mathrm{fs}$ 1 $\mathrm{fs}$ amber/protein.ff14SBonlysc implicit/gbn2
@@ B.2 Model configurations
We report the model configurations used to produce the results from PIPS [14] and TPS-GFN in Table 4 and Table 5, respectively.
Table 4: Model configurations used to produce the results from PIPS.
Molecule Rollout LR Samples Brownian motion std
Alanine Dipeptide 15000 $10^{-5}$ 16 0.1
Polyproline Helix 500 $3\times 10^{-5}$ 25 0.1
Chignolin 500 $10^{-4}$ 16 0.05
Table 5: Model configurations used to produce the results from TPS-GFN.
Molecule Force Flexible Rollout Trains per rollout LR Log Z LR Samples Target std($\sigma$)
Alanine Dipeptide $10000$ $2000$ $0.01$ $0.001$ $16$ $0.05$
Alanine Dipeptide $10000$ $2000$ $0.01$ $0.001$ $16$ $0.05$
Polyproline Helix $5000$ $200$ $0.0001$ $0.01$ $2$ $0.2$
Polyproline Helix $5000$ $200$ $0.001$ $0.01$ $2$ $0.2$
Chignolin $5000$ $200$ $0.0001$ $0.01$ $2$ $0.2$
Chignolin $5000$ $200$ $0.0001$ $0.01$ $2$ $0.2$
For Alanine Dipeptide, we remove the flexibility of length in our method by replacing our relaxed indicator function $\tilde{\mathbf{1}}_{\mathcal{B}}(\bm{x}_{1:T})$ with $\tilde{\mathbf{1}}_{\mathcal{B}}(\bm{x}_{T})=\exp\left(-\lVert D(\bm{r}_{T})-D(\bm{r}_{\mathcal{B}})\rVert_{F}^{2}/2\sigma^{2}\right)$ which slightly increases performance. However, other molecules benefit from flexiblity.
@@ B.3 Evaluation metrics
Following the prior work [14], we adopt three metrics for evaluation: (i) Expected pairwise distance (EPD), (ii) Target hit percentage (THP), and (iii) Energy transition point (ETP).
@@
The interatomic pairwise distance matrix $D(\cdot)$ [43] expresses molecular conformation in an E(3)-invariant manner. EPD measures the similarity between the final conformation $\bm{r}_{t}$ and the target conformation $\bm{r}_{\mathcal{B}}$ by calculating the square distance between interatomic pairwise distance matrices defined as follows:
$\operatorname{EPD}(\bm{r}_{t},\bm{r}_{\mathcal{B}})=\frac{1}{N^{2}}\lVert D(\bm{r}_{t})-D(\bm{r}_{\mathcal{B}})\rVert_{F}^{2}$
Where $t$ is the index of the final conformation of transition paths.
@@
While EPD continuously measures the distance from the fixed target conformation, THP measures the success of paths arriving at the target meta-stable state $\mathcal{B}$ in a binary manner. For the final conformations $\{r^{(i)}\}^{M}_{i=1}$ of $M$ paths, THP is defined as follows:
@@
For the alanine dipeptide, we use =$\mathcal{B}=\{r:|\phi(r)-\phi(r_{\mathcal{B}})|+|\psi(r)-\psi(r_{\mathcal{B}})|<0.75\}$ where $(\phi,\psi)$ is the collective variable consisting of two torsion angles.
@@
Unlike EPD and THP which evaluate the final conformation of a transition path, ETP measures the ability of the method to find the transition states when crossing the energy barrier. ETP refers to the maximum potential energy among states in a transition path. Formally, given a transition path $\bm{x}_{0:t}$ that reaches the target meta-stable region i.e., $\bm{r}_{t}\in\mathcal{B}$, ETP computes
$\operatorname{ETP}(\bm{x}_{0:t})=\max_{i\in[0,t]}U(\bm{r}_{i})$
However, we point out that these metrics can not measure the realisticity of the paths with respect to unbiased molecular dynamics. Thus, we propose metrics for the realisticity of paths and final conformations. The likelihood of the unbiased molecular dynamics is a part of the target distribution $p_{\mathcal{A},\mathcal{B}}(\bm{x}_{1:T}|\bm{x}_{0})$ which explains how realistic the sampled transition paths are.
@@
Log Likelihood (LL) refers to the log-likelihood of unbiased molecular dynamics as follows:
@@
We divide by $T$ to evaluate paths regardless of length. For the realisticity of final conformations, we focus on the potential energy of them. We propose two metrics, potential Energy of the Final Point (EFP) and Expected Pairwise Coulomb Distance (EPCD).
@@
The final conformation of a transition path from unbiased molecular dynamics should have low potential energy due to meta-stablity. EFP measures the potential $U(r_{t})$ of the final conformation of a transition path sampled from the model.
@@
Coulomb matrix [48] $C(r)$ of the final conformation $r$ describes electrostatic interaction between nuclei in the final conformation $r$ and is formulated as follows:
$C(r)_{ij}=\begin{cases}0.5Z_{i}^{2.4}&\text{for }i=j\\ Z_{i}Z_{j}/\lVert r_{i}-r_{j}\rVert&\text{for }i\neq j\end{cases}$
where $Z_{i}$ and $Z_{j}$ are the atomic numbers and $r_{i}$, $r_{j}$ are the positions of atoms $i$ and $j$, respectively. The atomic numbers $Z_{i}$ and $Z_{j}$ of the Coulomb matrix reflect atomic mass and energy level around the final conformation better than the distance matrix. We propose EPCD which replaces the distance metrics used in EPD with the Coulomb matrix and is defined as follows:
$\operatorname{EPCD}(\bm{r}_{t},\bm{r}_{\mathcal{B}})=\frac{1}{N^{2}}\lVert C(\bm{r}_{t})-C(\bm{r}_{\mathcal{B}})\rVert_{F}^{2}$
EPCD measures both geometric and electrostatic similarities between the final conformation $\bm{r}_{t}$ and the target conformation $\bm{r}_{\mathcal{B}}$.
Appendix C Additional visualizations
In this section, we provide additional visualizations of transition paths for each molecular system, generated by our TPS-GFN.
C.1 Alanine Dipeptide
In Figure 6, we visualize a collective variable (CV), i.e., two dihedral angles $(\phi,\psi)$ along the time horizon and 3D views of a transition path from the meta-stable region $C5$ to $C7_{ax}$ with the target conformation $\bm{r}_{C7_{ax}}$. Our model successfully generates a transition path since the final frame in Figure 6(l) is in $C7_{ax}$.
(a) Visualization of $\psi$ and $\phi$ along the time horizon
(b) Frame 0 $(=\bm{r}_{C5})$
(c) Frame 50
(d) Frame 100
(e) Frame 150
(f) Frame 200
(g) Frame 250
(h) Frame 300
(i) Frame 350
(j) Frame 400
(k) Frame 450
(l) Frame 498
(m) $\bm{r}_{C7_{ax}}$
Figure 6: 3D view of a transition path from the meta-stable region $C5$ to $C7_{ax}$ with the target conformation $\bm{r}_{C7_{ax}}$
C.2 Polyproline Helix
In Figure 7, we visualize 3D views of a transition path from the meta-stable region PP-II to PP-I.
(a) Frame 0 $(=\bm{r}_{\text{PP-II}})$
(b) Frame 40
(c) Frame 80
(d) Frame 120
(e) Frame 160
(f) Frame 200
(g) Frame 240
(h) Frame 280
(i) Frame 320
(j) Frame 357
(k) $\bm{r}_{\text{PP-I}}$
Figure 7: 3D view of a transition path from the meta-stable region PP-II to PP-I with the target conformation $\bm{r}_{\text{PP-I}}$
C.3 Chignolin
In Figure 8, we visualize 3D views of a Chignolin folding process.
(a) Frame 0 $(=\bm{r}_{\text{unfolded}})$
(b) Frame 100
(c) Frame 200
(d) Frame 300
(e) Frame 400
(f) Frame 500
(g) Frame 600
(h) Frame 700
(i) Frame 800
(j) Frame 900
(k) Frame 992
(l) $\bm{r}_{\text{folded}}$
Figure 8: 3D visualization of a Chignolin folding process with the target conformation $\bm{r}_{\text{folded}}$

### v3

@@ 4.3 Qualitative results
In this section, we evaluate our model based on visualizations of paths. For Alanine Dipeptide, we compare transition paths from our method with baselines for diversity. For Polyproline Helix and Chignolin, we visualize the 3D conformations of a transition path with a potential energy plot. For more visualizations, please refer to project page.
Alanine Dipeptide. In Figure 3, we sample transition paths of three methods from $C5$ to $C7_{ax}$ on Ramachandran plot of Alanine Dipeptide. We sampled 64 transition paths for all cases. The paths from unbiased MD simulations in Figure 3(a) remain trapped in the initial meta-stable region, and fail to produce transition paths. Both PIPS (Holdijk et al., 2024) and TPS-GFN generate valid transition paths between two meta-stable regions as shown in Figure 3(b) and Figure 3(c), respectively. PIPS generates transition paths that pass near one of the transition states while TPS-GFN generates transition paths that pass near both transition states as shown in Figure 3(c). This shows that TPS-GFN captures diverse transition paths and discovers transition states.
Polyproline Helix. We plot potential energy over time and visualize 3D structures of three states: initial, transition, and final state of a transition path. We also consider handedness, $H$, which differentiates between PP-II and PP-I (Moradi et al., 2009). Specifically, $H<0$ indicates a left-handed helix, while $H>0$ indicates a right-handed helix. In Figure 4, our method captures the isomerization of Polyproline Helix from left-handed to right-handed Helix.
Chignolin. Similarly, we plot potential energy and visualize 3D structures. We consider the donor-acceptance distance of the two key hydrogen bonds, Asp3OD–Thr6OG and Asp3N–Thr8O. In Figure 5, our method captures the folding process, by forming the key hydrogen bonds at the final state. Notably, our method generates folding processes about $4$ $\mathrm{ps}$, which is shorter than the average folding time of $0.6$ $\mathrm{\mu s}$ reported by Lindorff-Larsen et al. (2011).
@@ 5 Conclusion
In this work, we introduced a novel CV-free approach to sampling transition paths using generative flow networks (GFlowNets) to train neural bias potentials for efficient and diverse sampling. By reformulating the problem as amortized energy-based sampling and employing the one-step trajectory balance (TB) objective, our approach significantly improves the discovery of realistic transition paths compared to existing methods. Evaluations on proteins such as Alanine Dipeptide, Polyproline Helix, and Chignolin demonstrate superior accuracy and diversity in identifying transition states. We believe that our works expand to more complex systems, with significant implications for drug discovery and material design.
@@ Appendix A Generative Flow Networks
In this section, we briefly introduce the concept of GFlowNets and explain how we derive Equation 2. We refer to Lahlou et al. (Lahlou et al., 2023) for a comprehensive study on formalizing GFlowNets on continuous state spaces.
Generative flow networks. GFlowNet (Bengio et al., 2021, 2023) is a framework to sequentially construct samples $s\in\mathcal{S}$ proportional to a reward function $R(s)$, where $\mathcal{S}$ is a set of data. In other words, GFlowNet aims to learn a target distribution $p^{\ast}(s)$ defined as:
@@
Trajectory balance. For flow matching, Malkin et al. (Malkin et al., 2022a) proposed the trajectory balance objective:
@@
where $P_{B}(s_{0}|s_{1};\theta)$ is ignored since $s_{0}$ is the constant initial state. Note that the detailed balance (Bengio et al., 2023) objective also reduces to the same formula as in Equation 4 with one-step GFlowNets.
@@ Appendix B Experiment details
We implement codes based on OpenMM (Eastman et al., 2023) and Pytorch (Paszke et al., 2019) libraries and run codes on a single GPU (NVIDIA RTX 3090 or RTX A5000).
@@ B.1 OpenMM configurations
We use the Langevin integrator (Bussi & Parrinello, 2007) with the step size $\Delta=1~{}\mathrm{fs}$ and the friction term $\gamma=1~{}\mathrm{ps}^{-1}$. In training TPS-GFN, we start simulations from a temperature $\lambda_{\text{start}}=600$K, linearly decrease temperature, and end at a temperature $\lambda_{\text{end}}=300$K. Other OpenMM configurations are shown in Table 3. For PIPS, the number of steps $T$ of Alanine Dipeptide is $500$ as given by Holdijk et al. (2024).
Table 3: OpenMM configurations.
Molecule ##\## steps ($T$) Force field Solvent
Alanine Dipeptide $1000$ amber99sbildn tip3p
Polyproline Helix $10000$ amber/protein.ff14SBonlysc implicit/gbn2
Chignolin $10000$ amber/protein.ff14SBonlysc implicit/gbn2
@@ B.2 Model configurations
For PIPS, we use the model configurations reported by Holdijk et al. (2024). For TPS-GFN, we roll out $I=1000$ times, set the scaling factor $h=8$, the learning rate to $0.0001$ for the neural bias potential, and $0.01$ for a scalar parameter $w=\log Z_{\theta}$. We concatenate the distance feature $d_{i}=\lVert\bm{r}_{i}-(\bm{r}_{\mathcal{B}})_{i}\rVert_{2}$ and the position $\bm{r}_{i}$ of the $i$-th atom, using this as input for the neural bias potential. We report other model configurations in Table 4. We use the same configurations for both the neural bias potential and force, except for the output dimension.
Table 4: Model configurations of TPS-GFN.
Molecule Trains per rollout ($J$) # Samples ($M_{1}$) Batch Size ($M_{2}$) Buffer Size Relaxation ($\sigma$)
Alanine Dipeptide $2000$ $16$ $16$ $2048$ $0.05$
Polyproline Helix $100$ $2$ $2$ $256$ $0.1$
Chignolin $100$ $2$ $2$ $256$ $0.2$
@@ B.3 Evaluation metrics
Following Holdijk et al. (2024), we adopt the following three metrics for evaluation.
@@
The interatomic pairwise distance matrix $D(\cdot)$ expresses molecular conformation in an E(3)-invariant manner. EPD measures the similarity between the final conformation $\bm{r}_{t_{\ast}}$ and the target conformation $\bm{r}_{\mathcal{B}}$ by comparing interatomic pairwise distance matrices as follows:
$\operatorname{EPD}(\bm{r}_{t_{\ast}},\bm{r}_{\mathcal{B}})=\frac{1}{N^{2}}\lVert D(\bm{r}_{t_{\ast}})-D(\bm{r}_{\mathcal{B}})\rVert_{F}^{2}$
Where $t_{\ast}$ is the index of the final conformation of transition paths.
@@
THP measures the success rate of paths arriving at the target meta-stable state $\mathcal{B}$ in a binary manner. For the final conformations $\{r^{(i)}\}^{M}_{i=1}$ of $M$ paths, THP is defined as follows:
@@
For the alanine dipeptide, we use =$\mathcal{B}=\{r:|\phi(r)-\phi(r_{\mathcal{B}})|+|\psi(r)-\psi(r_{\mathcal{B}})|<0.75\}$ where $(\phi,\psi)$ is the collective variable consisting of two backbone dihedral angles.
@@
ETP measures the ability of the method to find the transition states when crossing the energy barrier. ETP refers to the maximum potential energy among states in a transition path. Formally, given a transition path $\bm{x}_{0:t_{\ast}}$ that reaches the target meta-stable region i.e., $\bm{r}_{t_{\ast}}\in\mathcal{B}$, ETP is defined as follows:
$\operatorname{ETP}(\bm{x}_{0:t_{\ast}})=\max_{i\in[0,t_{\ast}]}U(\bm{r}_{i})$
However, we point out that these metrics can not measure the realisticity of the overall paths with respect to unbiased molecular dynamics (MD) and final conformations with respect to their potential energy.
@@
Log-likelihood (LL) evaluates the realisticity of paths with respect to unbiased MD simulation at 300K, referring to the log-likelihood of conditional density of unbiased MD simulation as follows:
@@
We divide by $T$ to evaluate paths regardless of length.
For the realisticity of final conformations, we focus on the potential energy of them. We propose two metrics, potential Energy of the final point (EFP) and Expected pairwise coulomb distance (EPCD).
@@
The final conformations of a transition path from unbiased MD simulation have low potential energy due to meta-stability. EFP measures the potential energy $U(r_{t_{\ast}})$ of the final conformations of transition paths sampled from the model.
@@
Coulomb matrix (Rupp et al., 2012) $C(r)$ of the conformation $r$ describes electrostatic interaction between nuclei and is formulated as follows:
$C(\bm{r})_{ij}=\begin{cases}0.5Z_{i}^{2.4}&\text{for }i=j\\ Z_{i}Z_{j}/\lVert\bm{r}_{i}-\bm{r}_{j}\rVert_{2}&\text{for }i\neq j\end{cases}$
where $Z_{i}$ and $Z_{j}$ are the atomic numbers and $r_{i}$, $r_{j}$ are the positions of atoms $i$ and $j$, respectively. The atomic numbers $Z_{i}$ and $Z_{j}$ of the Coulomb matrix reflect mass and energy level around the final conformation better than the distance matrix. We propose EPCD which replaces the distance metrics used in EPD with the Coulomb matrix and is defined as follows:
$\operatorname{EPCD}(\bm{r}_{t},\bm{r}_{\mathcal{B}})=\frac{1}{N^{2}}\lVert C(\bm{r}_{t_{\ast}})-C(\bm{r}_{\mathcal{B}})\rVert_{F}^{2}$
EPCD measures geometric and electrostatic similarities between the final conformation $\bm{r}_{t_{\ast}}$ and the target conformation $\bm{r}_{\mathcal{B}}$.